If the sides of a square are increased by 3 in., the area is increased by
39 in.2. What were the dimensions of the original square?

Answer by ankor@dixie-net.com(22740)

If the sides of a square are increased by 3 in., the area is increased by
39 in.2. What were the dimensions of the original square?
:
Let x = the side of the original square
Then
x^2 = the area of the original square
and
(x+3) = the side of the new square
then
(x+3)^2 = the area of the new square
:
New square area - old square area = 39
(x+3)^2 - x^2 = 39
:
FOIL (x+2)(x+3)
x^2 + 6x + 9 - x^2 = 39
:
Conveniently the x^2's eliminate each other so we have:
6x + 9 = 39
6x = 39-9
x = 30/6
x = 5 in, side of the original square
:
:
Check our solution, (new square side would be 8")
8^2 - 5^2
64 - 25 = 39 sq/in as given
:
Did you understand what went on here? Any questions?